# Finding all solutions of qKZ equations in characteristic $p$

### v1

@@
Abstract.
In [MV] the difference qKZ equations were considered modulo a prime number $p$ and a family of polynomial solutions of the qKZ equations modulo $p$ was constructed by an elementary procedure as suitable $p$-approximations of the hypergeometric integrals. In this paper, we study in detail the first family of nontrivial example of the qKZ equations in characteristic $p$. We describe all solutions of these qKZ equations in characteristic $p$ by demonstrating that they all stem from the $p$-hypergeometric solutions. We also prove a Lagrangian property (called the orthogonality property) of the subbundle of the qKZ bundle spanned by the $p$-hypergeometric sections. This paper extends the results of [VV1] on the differential KZ equations to the difference qKZ equations.
$\kern-3.99994pt^{\circ}\mskip-0.99998mu$epartment of Mathematical Sciences, Indiana University Indianapolis
402 North Blackford St, Indianapolis, IN 46202-3216, USA
Department of Mathematics, University of North Carolina at Chapel Hill
Chapel Hill, NC 27599-3250, USA
Key words: qKZ equations, $p$-hypergeometric solutions, orthogonality relations, $p$-curvature
2020 Mathematics Subject Classification: 11D79 (12H25, 32G34, 33C05, 33E30)
footnotetext: ${}^{\circ}\mskip-0.99998mu$-mail: emukhin@iupui.edu, supported in part by Simons Foundation grant 709444
${}^{\star}\mskip-0.99998mu$-mail: anv@email.unc.edu
1. Introduction
The Knizhnik-Zamolodchikov (KZ) differential equations are a system of linear differential equations, satisfied by conformal blocks on the sphere in the WZW model of conformal field theory, see [KZ]. The quantum Knizhnik-Zamolodchikov (qKZ) equations are a difference version of the KZ equations which naturally appear in the representation theory of Yangians (rational case) and quantum affine algebras (trigonometric case), see [FR, EFK]. The qKZ equations are a deformation of the KZ differential equations.
@@ 1. Introduction
In [SV] the differential KZ equations were considered modulo a prime integer $p$. It turned out that modulo $p$ the KZ equations have a family of polynomial solutions. The construction of these solutions was analogous to the construction of the multidimensional hypergeometric solutions, and these polynomial solutions were called the $p$-hypergeometric solutions.
In [MV] the rational ${\mathfrak{sl}_{2}}$ qKZ equations with values in the $n$-th tensor power of the vector representation $L$ and an integer step $\kappa$ were considered modulo $p$. A family of polynomial solutions modulo $p$ of these equations was constructed and called the $p$-hypergeometric solutions.
In this paper we address the problem of whether all solutions of the qKZ equations in characteristic $p$ are generated by the $p$-hypergeometric solutions. We consider the first family of nontrivial examples of the qKZ equations and demonstrate that, indeed, in this case, all solutions of the qKZ equations stem from the $p$-hypergeometric solutions.
@@
where the linear operators $K_{a}(z;\kappa)$ are given in terms of the rational ${\mathfrak{sl}_{2}}$ $R$-matrix, see (2.3). The operators $K_{a}(z;\kappa)$ commute with the diagonal action of ${\mathfrak{sl}_{2}}$, and, therefore, it is sufficient to solve the qKZ equations only with values in the space of singular vectors of a given weight. In this paper we study the qKZ equations with values in $V:={\operatorname{Sing}}L^{\otimes n}[n-2]\subset L^{\otimes n}$, the subspace of singular vectors of weight $n-2$. We have $\dim V=n-1$.
@@
Assume that $\kappa\in{\mathbb{F}}_{p}^{\times}$. Let $0<k<p$ be the positive integer such that $\tilde{\kappa}k\,\equiv\,-1\pmod{p}$. Let $[x]$ denote the integer part of a real number and $d(\kappa):=\left[\frac{kn}{p}\right]$. If $p$ does not divide $n$, then $d(\kappa)+d(-\kappa)=n-1=\dim V$.
In [MV] we constructed $d(\kappa)$ $V$-valued $p$-hypergeometric solutions of the qKZ equations denoted $Q^{\ell p-1}(z;\kappa)$, $\ell=1,\dots,d(\kappa)$. In this paper we show that these solutions are linearly independent over the field ${\mathbb{K}}(z_{1},\dots,z_{n})$, see Theorem 5.6.
@@
Let $p>n$, $\kappa\in{\mathbb{F}}_{p}^{\times}$, and $10<d(\kappa)<n-$ 1. Let $f(z)$ be a $V$-valued rational function in $z$ which is a solution the qKZ equations with step $\kappa$. Then $f(z)$ is a linear combination of the $p$-hypergeometric solutions $Q^{\ell p-1}(z;\kappa)$, $\ell=1,\dots,d(\kappa)$, with coefficients which are scalar rational functions in $z_{i}^{p}-z_{i}$, $i=1,\dots,n$.
@@
If $d(\kappa)=n-1$ or 00, all solutions of the qKZ equations with values in $V$ and step $\kappa$ are described in Section 7.4.
We prove the orthogonality relations for $p$-hypergeometric solutions of the qKZ equations with steps $\kappa$ and $-\kappa$.
@@
Let $p>n$ and $10<d(\kappa)<n-$ 1. Then for any $\ell\in\{1,\dots,d(\kappa)\}$ and $m\in\{1,\dots,d(-\kappa)\}$, we have
(1.1) $S\big{(}Q^{mp-1}(-z;-\kappa),Q^{\ell p-1}(z;\kappa)\big{)}=0,$
@@
$C_{a}(z_{1},\dots,z_{n};\kappa)$ ::=:= $K_{a}(z_{1},\dots,z_{a}-(p-1)\kappa,\dots,z_{n};\kappa)\times$
$\times$ $K_{a}(z_{1},\dots,z_{a}-(p-2)\kappa,\dots,z_{n};\kappa)\dots K_{a}(z_{1},\dots,z_{a}-\kappa,\dots,z_{n};\kappa)\times$
$\times$ $K_{a}(z_{1},\dots,z_{a},\dots,z_{n};\kappa),$
@@ 2. $\operatorname{(ii)}$
    If $\kappa\in{\mathbb{F}}_{p}^{\times}$ and $d(\kappa)=n-1$ or 00, then all reduced $p$-curvature operators $\hat{C}_{a}(z;\kappa)$, $a=1,\dots,n$, are equal to zero.
@@ 3. $\operatorname{(iii)}$
    If $\kappa\in{\mathbb{F}}_{p}^{\times}$ and $10<d(\kappa)<n-$ 1, then all reduced $p$-curvature operators $\hat{C}_{a}(z;\kappa)$, $a=1,\dots,n$, are nonzero. For every $a$, the span of the $p$-hypergeometric solutions $Q^{\ell p-1}(z;\kappa)$, $\ell=1,\dots,d(\kappa)$, lies in the kernel of $\hat{C}_{a}(z,\kappa)$ and contains the image of $\hat{C}_{a}(z;\kappa)$. Also, for all $a,b$,
    $\hat{C}_{a}(z;\kappa)\hat{C}_{b}(z;\kappa)$ $=$ $0,$
    $\hat{C}_{a}(z,-\kappa)+\hat{C}_{a}(-z;\kappa)^{*}$ $=$ 00 0
    where for an operator $T:V\to V$ we denote by $T^{*}$ the operator dual to $T$ under the Shapovalov form.
@@
In [VV1] the differential KZ equations over a field ${\mathbb{K}}$ of characteristic $p$ with values in $V\subset L^{\otimes n}$ were studied in detail. Our paper extends the results of [VV1] from the differential KZ equations to the difference qKZ equations. The proofs of Theorems 1.1, 1.2, and 1.4 are based on the corresponding results in [VV1] for the differential KZ equations.
On the differential and difference equations in characteristic $p$ and associated $p$-curvature see also [EV, VV2].
Acknowledgments
The authors thank P. Etingof, R. Rimanyi, V. Vologodsky for useful discussions.
2. Difference qKZ equations
2.1. Notations
In these paper $p$ is a prime and ${\mathbb{K}}$ a field of characteristic $p$.
Consider the Lie algebra ${\mathfrak{sl}_{2}}$ over ${\mathbb{K}}$ with basis $e,f,h$ and relations $[e,f]=h$, $[h,e]=2e$, $[h,f]=-2f$. Let $L$ be the two-dimensional ${\mathfrak{sl}_{2}}$-module with basis $v_{1},v_{2}$ and the action $ev_{1}=0,ev_{2}=v_{1}$, $fv_{1}=v_{2},fv_{2}=0$, $hv_{1}=v_{1},hv_{2}=-v_{2}$.
@@ 2.1. Notations
where $i_{j}=2$ if $i_{j}\in I$ and $i_{j}=1$ if $i_{j}\notin I$. Denote by $L^{\otimes n}[n-2l]$ the span of the vectors $\{v_{I}\ |\ I\in\mathcal{I}_{l}\}$. We have a direct decomposition,
$L^{\otimes n}=\oplus_{l=0}^{n}L^{\otimes n}[n-2l].$
@@
2.2. qKZ equations
Define the rational $\,R$-matrix acting on $\>L^{\otimes 2}$,
$R(u)\,=\,\frac{u-P}{u-1}\,,\ignorespaces$
where $P$ is the permutation of tensor factors of $L^{\otimes 2}$. The $R$-matrix satisfies the Yang-Baxter and unitarity equations,
(2.1) $R^{(12)}(u-v)R^{(13)}(u)R^{(23)}(v)\,=\,R^{(23)}(v)R^{(13)}(u)R^{(12)}(u-v)\,,$
@@
The first equation is an equation in $\operatorname{End}(L^{\otimes 3})$. The superscript indicates the factors of $\>L^{\otimes 3}$ on which the corresponding operators act.
Let $\,z=(z_{1},\allowbreak\dots,\allowbreak z_{n})$. Define the qKZ operators $\,K_{1},\dots,K_{n}\mskip 2.00002mu$ acting on $L^{\otimes n}$:
(2.3) $K_{a}(z;\kappa)=R^{(a,a-1)}(z_{a}-z_{a-1}-\kappa)\,\dots\,R^{(a,1)}(z_{a}-z_{1}-\kappa)$
×$\hskip 120.0pt\times\ R^{(a,n)}(z_{a}-z_{n})\,\dots\,R^{(a,a+1)}(z_{a}\mskip-0.99998mu-z_{a+1})\,,$
@@
$K_{a}(z_{1},\dots,z_{b}-\kappa,\dots,z_{n};\kappa)\,K_{b}(z;\kappa)=K_{b}(z_{1},\dots,z_{a}-\kappa\dots,z_{n};\kappa)\,K_{a}(z;\kappa)$
for $a,b=1,\dots,n$, see [FR].
@@
(2.4) $s(z_{1},\dots,z_{a}-\kappa,\dots,z_{n})\,=\,K_{a}(z\>;\kappa)\,s(z),\qquad a=1,\dots,n,\ignorespaces$
@@
(2.5) $({\bf T}_{a}f)(z_{1},\dots,z_{a},\dots,z_{n})=f(z_{1},\dots,z_{a}-\kappa,\dots,z_{n}).$
The difference operators
$\nabla_{a}={\bf T}_{a}^{-1}K_{a}(z;\kappa)$
are called the connection operators of the qKZ difference connection. We have $[\nabla_{a},\nabla_{b}]=0$.
2.3. $p$-curvature of the qKZ connection
@@
$C_{a}(z;\kappa)$ $=$ $K_{a}(z_{1},\dots,z_{a}-(p-1)\kappa,\dots,z_{n};\kappa)K_{a}(z_{1},\dots,z_{a}-(p-2)\kappa,\dots,z_{n};\kappa)\dots$
$\dots\ K_{a}(z_{1},\dots,z_{a}-\kappa,\dots,z_{n};\kappa)K_{a}(z_{1},\dots,z_{a},\dots,z_{n};\kappa),$
for $a=1,\dots,n$. In other words,
(2.7) $C_{a}=(\nabla_{a})^{p}\,.$
@@
The operators $C_{a}(z;\kappa)$ commute, $[C_{a}(z;\kappa),C_{b}(z;\kappa)]=0$.
@@
then $s(z;\kappa)$ is an eigenvector of the the $p$-curvature operators with eigenvalue 1,
(2.8) $s(z;\kappa)=C_{a}(z;\kappa)s(z;\kappa),\qquad a=1,\dots,n.$
@@
$=$ $\prod_{j\neq a}(z_{a}^{p}-\kappa^{p-1}z_{a}+(-z_{j})^{p}+\kappa^{p-1}z_{j}+(-1)^{p}+\kappa^{p-1}).$
@@
(2.10) $\tilde{C}_{a}(z;\kappa)=D_{a}(z;\kappa)(C_{a}(z;\kappa)-1).$
@@ 2.4. Differential KZ equations
(2.11) $\nabla_{a}^{\operatorname{KZ}}=\kappa\,\frac{\partial}{\partial z_{a}}-\sum_{j\neq a}\frac{P^{(a,j)}-1}{z_{a}-z_{j}}\,,\qquad a=1,\dots,n,$
define a flat KZ connection on $L^{\otimes n}\times K^{n}\to{\mathbb{K}}^{n}$, $[\nabla_{a}^{\operatorname{KZ}},\nabla_{b}^{\operatorname{KZ}}]=0$. The operators $\nabla_{a}^{\operatorname{KZ}}$ commute with the ${\mathfrak{sl}_{2}}$-action on $L^{\otimes n}$. The system of equations
(2.12) $\kappa\,\frac{\partial f}{\partial z_{a}}=\sum_{j\neq a}\frac{P^{(a,j)}-1}{z_{a}-z_{j}}\,f,\qquad a=1,\dots,n,$
@@
(2.14) $C_{a}^{\operatorname{KZ}}(z;\kappa):=\left(\nabla_{a}^{\operatorname{KZ}}\right)^{p}\,.$
They define an endomorphism of the KZ connection, $[C_{a}^{\operatorname{KZ}},\nabla_{b}^{\operatorname{KZ}}]=0$.
@@
(2.15) $\phantom{aaa}D^{\operatorname{KZ}}_{a}(z;\kappa)=\prod_{j\neq a}(z_{a}^{p}-z_{j}^{p}).$
@@
(2.16) $\tilde{C}^{\operatorname{KZ}}_{a}(z;\kappa)=D^{\operatorname{KZ}}_{a}(z;\kappa)C_{a}^{\operatorname{KZ}}(z;\kappa)\,.$
The normalized $p$-curvature operator is a homogeneous polynomial in $z$ of degree $(n-2)p$ if nonzero.
2.5. KZ equations as a limit of qKZ equations

### v2

@@
[…]
Finding All Solutions of qKZ Equations in Characteristic $p$
[…]
Finding All Solutions of qKZ Equations
in Characteristic $\boldsymbol{p}$
[…]
Evgeny MUKHIN a and Alexander VARCHENKO b
[…]
E. Mukhin and A. Varchenko
[…]
a) Department of Mathematical Sciences, Indiana University Indianapolis,
[…]
b) Department of Mathematics, University of North Carolina at Chapel Hill,
[…]
Received September 23, 2025, in final form December 16, 2025; Published online January 02, 2026
[…]
In [J. Lond. Math. Soc. 109 (2024), e12884, 22 pages], the difference qKZ equations were considered modulo a prime number $p$ and a family of polynomial solutions of the qKZ equations modulo $p$ was constructed by an elementary procedure as suitable $p$-approximations of the hypergeometric integrals. In this paper, we study in detail the first family of nontrivial examples of the qKZ equations in characteristic $p$. We describe all solutions of these qKZ equations in characteristic $p$ by demonstrating that they all stem from the $p$-hypergeometric solutions. We also prove a Lagrangian property (called the orthogonality property) of the subbundle of the qKZ bundle spanned by the $p$-hypergeometric sections. This paper extends the results of [arXiv:2405.05159] on the differential KZ equations to the difference qKZ equations.
[…]
qKZ equations; $p$-hypergeometric solutions; orthogonality relations; $p$-curvature
[…]
11D79; 12H25; 32G34; 33C05; 33E30
1 Introduction
The Knizhnik–Zamolodchikov (KZ) differential equations are a system of linear differential equations, satisfied by conformal blocks on the sphere in the WZW model of conformal field theory, see [4]. The quantum Knizhnik–Zamolodchikov (qKZ) equations are a difference version of the KZ equations which naturally appear in the representation theory of Yangians (rational case) and quantum affine algebras (trigonometric case), see [1, 3]. The qKZ equations may be regarded as a deformation of the KZ differential equations.
@@ 1 Introduction
In [6], the differential KZ equations were considered modulo a prime integer $p$. It turned out that modulo $p$ the KZ equations have a family of polynomial solutions. The construction of these solutions was analogous to the construction of the multidimensional hypergeometric solutions, and these polynomial solutions were called the $p$-hypergeometric solutions.
In [5], the rational $\mathfrak{sl}_{2}$} qKZ equations with values in the $n$-th tensor power of the vector representation $L$ and an integer step $\kappa$ were considered modulo $p$. A family of polynomial solutions modulo $p$ of these equations was constructed and called the $p$-hypergeometric solutions.
In this paper, we address the problem of whether all solutions of the qKZ equations in characteristic $p$ are generated by the $p$-hypergeometric solutions. We consider the first family of nontrivial examples of the qKZ equations and demonstrate that, indeed, in this case, all solutions of the qKZ equations stem from the $p$-hypergeometric solutions.
@@
where the linear operators $K_{a}(z;\kappa)$ are given in terms of the rational $\mathfrak{sl}_{2}$} $R$-matrix, see (2.2). The operators $K_{a}(z;\kappa)$ commute with the diagonal action of $\mathfrak{sl}_{2}$}, and, therefore, it is sufficient to solve the qKZ equations only with values in the space of singular vectors of a given weight. In this paper, we study the qKZ equations with values in $V:=\operatorname{Sing}L^{\otimes n}[n-2]\subset L^{\otimes n}$, the subspace of singular vectors of weight n2n-2. We have $\dim V=n-1$.
@@
Assume that $\kappa\in{\mathbb{F}}_{p}^{\times}$. Let $0<k<p$ be the positive integer such that $\kappa k\equiv-1\pmod{p}$}. Let $[x]$ denote the integer part of a real number and $d(\kappa):=\big[\frac{kn}{p}\big]$. If $p$ does not divide $n$, then $d(\kappa)+d(-\kappa)=n-1=\dim V$}.
In [5], we constructed $d(\kappa)$ $V$-valued $p$-hypergeometric solutions of the qKZ equations denoted $Q^{\ell p-1}(z;\kappa)$}, $\ell=1,\dots,d(\kappa)$. In this paper, we show that these solutions are linearly independent over the field ${\mathbb{K}}(z_{1},\dots,z_{n})$, see Theorem 5.6.
@@
Let $p>n$, $\kappa\in{\mathbb{F}}_{p}^{\times}$, and $0<d(\kappa)<n-1$. Let $f(z)$ be a $V$-valued rational function in $z$ which is a solution the qKZ equations with step $\kappa$. Then $f(z)$ is a linear combination of the $p$-hypergeometric solutions $Q^{\ell p-1}(z;\kappa)$, $\ell=1,\dots,d(\kappa)$, with coefficients which are scalar rational functions in $z_{i}^{p}-z_{i}$, $i=1,\dots,n$.
@@
If $d(\kappa)=n-1$ or 0, all solutions of the qKZ equations with values in $V$ and step $\kappa$ are described in Section 7.4.
We prove the orthogonality relations for $p$-hypergeometric solutions of the qKZ equations with steps $\kappa$ and $\kappa$.
@@
Let $p>n$ and $0<d(\kappa)<n-1$. Then for any $\ell\in\{1,\dots,d(\kappa)\}$ and $m\in\{1,\dots,d(-\kappa)\}$, we have
$S\big(Q^{mp-1}(-z;-\kappa),Q^{\ell p-1}(z;\kappa)\big)=0,$
@@
$C_{a}(z_{1},\dots,z_{n};\kappa):=K_{a}(z_{1},\dots,z_{a}-(p-1)\kappa,\dots,z_{n};\kappa)$
[…]
@@ 2. $\operatorname{(ii)}$
    If $\kappa\in{\mathbb{F}}_{p}^{\times}$ and $d(\kappa)=n-1$ or 0, then all reduced $p$-curvature operators $\hat{C}_{a}(z;\kappa)$, $a=1,\dots,n$, are equal to zero.
@@ 3. $\operatorname{(iii)}$
    If $\kappa\in{\mathbb{F}}_{p}^{\times}$ and $0<d(\kappa)<n-1$, then all reduced $p$-curvature operators $\hat{C}_{a}(z;\kappa)$, $a=1,\dots,n$, are nonzero. For every $a$, the span of the $p$-hypergeometric solutions $Q^{\ell p-1}(z;\kappa)$, $\ell=1,\dots,d(\kappa)$, lies in the kernel of $\hat{C}_{a}(z,\kappa)$ and contains the image of $\hat{C}_{a}(z;\kappa)$. Also, for all $a$, $b$,
    $\hat{C}_{a}(z;\kappa)\hat{C}_{b}(z;\kappa)=0,\qquad\hat{C}_{a}(z,-\kappa)+\hat{C}_{a}(-z;\kappa)^{*}=0,$
    where for an operator $T\colon V\to V$ we denote by $T^{*}$ the operator dual to $T$ under the Shapovalov form.
@@
In [8], the differential KZ equations over a field $\mathbb{K}$} of characteristic $p$ with values in $V\subset L^{\otimes n}$ were studied in detail. Our paper extends the results of [8] from the differential KZ equations to the difference qKZ equations. The proofs of Theorems 1.1, 1.2, and 1.4 are based on the corresponding results in [8] for the differential KZ equations.
On the differential and difference equations in characteristic $p$ and associated $p$-curvature see also [2, 9].
2 Difference qKZ equations
2.1 Notations
In this paper, $p$ is a prime and $\mathbb{K}$} a field of characteristic $p$.
Consider the Lie algebra $\mathfrak{sl}_{2}$} over $\mathbb{K}$} with basis $e$, $f$, $h$ and relations $[e,f]=h$, $[h,e]=2e$, [h,f]=2f[h,f]=-2f. Let $L$ be the two-dimensional $\mathfrak{sl}_{2}$}-module with basis $v_{1}$, $v_{2}$ and the action $ev_{1}=0$, $ev_{2}=v_{1}$, $fv_{1}=v_{2}$, $fv_{2}=0$, $hv_{1}=v_{1}$, $hv_{2}=-v_{2}$.
@@ 2.1 Notations
where $i_{j}=2$ if $i_{j}\in I$ and $i_{j}=1$ if $i_{j}\notin I$. Denote by $L^{\otimes n}[n-2l]$ the span of the vectors $\{v_{I}\mid I\in\mathcal{I}_{l}\}$. We have a direct sum decomposition,
$L^{\otimes n}=\bigoplus_{l=0}^{n}L^{\otimes n}[n-2l].$
@@
2.2 qKZ equations
Define the rational $R$-matrix acting on $L^{\otimes 2}$, $R(u)=\frac{u-P}{u-1}$, where $P$ is the permutation of tensor factors of $L^{\otimes 2}$. The $R$-matrix satisfies the Yang–Baxter and unitarity equations,
$R^{(12)}(u-v)R^{(13)}(u)R^{(23)}(v)=R^{(23)}(v)R^{(13)}(u)R^{(12)}(u-v),$
$R^{(12)}(u)R^{(21)}(-u)=1.$ (2.1)
The first equation is an equation in $\operatorname{End}\big(L^{\otimes 3}\big)$. The superscript indicates the factors of $L^{\otimes 3}$ on which the corresponding operators act.
Let $z=(z_{1},\allowbreak\dots,\allowbreak z_{n})$. Define the qKZ operators $K_{1},\dots,K_{n}\mskip 2.00002mu$ acting on $L^{\otimes n}$
$K_{a}(z;\kappa)=R^{(a,a-1)}(z_{a}-z_{a-1}-\kappa)\cdots R^{(a,1)}(z_{a}-z_{1}-\kappa)$
@@
$K_{a}(z_{1},\dots,z_{b}-\kappa,\dots,z_{n};\kappa)K_{b}(z;\kappa)=K_{b}(z_{1},\dots,z_{a}-\kappa,\dots,z_{n};\kappa)K_{a}(z;\kappa)$
for $a,b=1,\dots,n$, see [3].
@@
$s(z_{1},\dots,z_{a}-\kappa,\dots,z_{n})=K_{a}(z;\kappa)s(z),\qquad a=1,\dots,n,$ (2.3)
@@
$({\bf T}_{a}f)(z_{1},\dots,z_{a},\dots,z_{n})=f(z_{1},\dots,z_{a}-\kappa,\dots,z_{n}).$
The difference operators $\nabla_{a}={\bf T}_{a}^{-1}K_{a}(z;\kappa)$ are called the connection operators of the qKZ difference connection. We have $\nabla_{a},\nabla_{b}]=0$.
2.3 $\boldsymbol{p}$-curvature of the qKZ connection
@@
$C_{a}(z;\kappa)=K_{a}(z_{1},\dots,z_{a}-(p-1)\kappa,\dots,z_{n};\kappa)K_{a}(z_{1},\dots,z_{a}-(p-2)\kappa,\dots,z_{n};\kappa)\cdots$
[…]
for $a=1,\dots,n$. In other words, $C_{a}=(\nabla_{a})^{p}$.
@@
The operators $C_{a}(z;\kappa)$ commute, i.e., $C_{a}(z;\kappa),C_{b}(z;\kappa)]=0$.
@@
then $s(z;\kappa)$ is an eigenvector of the $p$-curvature operators with eigenvalue 1,
$s(z;\kappa)=C_{a}(z;\kappa)s(z;\kappa),\qquad a=1,\dots,n.$ (2.4)
@@
$={}$ $\prod_{j\neq a}\big(z_{a}^{p}-\kappa^{p-1}z_{a}+(-z_{j})^{p}+\kappa^{p-1}z_{j}+(-1)^{p}+\kappa^{p-1}\big).$
@@
$\tilde{C}_{a}(z;\kappa)=D_{a}(z;\kappa)(C_{a}(z;\kappa)-1).$
@@ 2.4 Differential KZ equations
$\nabla_{a}^{\operatorname{KZ}}=\kappa\frac{\partial}{\partial z_{a}}-\sum_{j\neq a}\frac{P^{(a,j)}-1}{z_{a}-z_{j}},\qquad a=1,\dots,n,$
define a flat KZ connection on $L^{\otimes n}\times K^{n}\to{\mathbb{K}}^{n}$, $\big[\nabla_{a}^{\operatorname{KZ}},\nabla_{b}^{\operatorname{KZ}}\big]=0$. The operators $\nabla_{a}^{\operatorname{KZ}}$ commute with the $\mathfrak{sl}_{2}$}-action on $L^{\otimes n}$. The system of equations
$\kappa\frac{\partial f}{\partial z_{a}}=\sum_{j\neq a}\frac{P^{(a,j)}-1}{z_{a}-z_{j}}f,\qquad a=1,\dots,n,$ (2.6)
@@
$C_{a}^{\operatorname{KZ}}(z;\kappa):=\left(\nabla_{a}^{\operatorname{KZ}}\right)^{p}.$
They define an endomorphism of the KZ connection, $\big[C_{a}^{\operatorname{KZ}},\nabla_{b}^{\operatorname{KZ}}\big]=0$.
@@
$D^{\operatorname{KZ}}_{a}(z;\kappa)=\prod_{j\neq a}\big(z_{a}^{p}-z_{j}^{p}\big).$
@@
$\tilde{C}^{\operatorname{KZ}}_{a}(z;\kappa)=D^{\operatorname{KZ}}_{a}(z;\kappa)C_{a}^{\operatorname{KZ}}(z;\kappa).$
The normalized $p$-curvature operator is a homogeneous polynomial in $z$ of degree (n2)p(n-2)p if nonzero.
2.5 KZ equations as a limit of qKZ equations